OpenStudy (anonymous):

Someone please help me with sine, cosine, tangent, etc.? Easy questions just need help

OpenStudy (anonymous):

@dpasingh please

OpenStudy (anonymous):

@washcaps Post your questions...

OpenStudy (anonymous):

I have to fill in a table that has empty cells. The table has cos(theta) tan(theta) csc(theta) sec(theta) cot(theta) sin(theta), theta in degrees, and theta in radians. The theta is the zero with a line crossing through it. Anyway, I have cos(theta) as 1 and I have to find the other values

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

@washcaps Post your questions..

OpenStudy (anonymous):

That is my question. I need to know how to fill in the other cells given cos(theta) 1

OpenStudy (anonymous):

Ok you know that \[\cos \theta = 1 \rightarrow \cos \theta = \cos 0^0 \rightarrow \theta = 0^0\] since \[\cos 0^0 =1 ( chek -from-the- table)\] Now we can find other trigonometric ratios, like \[\sin \theta = \sin 0^0 = 0\] \[\tan \theta = \tan0^0 = \infty \] \[\cot \theta = \cot 0^0 = 0\] \[\sec \theta = \sec 0^0 = 1\] \[cosec \theta = cosec 0^0 = \infty\] @washcaps

OpenStudy (anonymous):

So those are the values I have to put into the empty cells? What is theta in degrees and theta in radians?

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how do I know the theta in degrees and theta in radians @dpasingh ?

OpenStudy (anonymous):

let \[\theta \] be angle in degrees so we write it as \[ \huge \theta ^0 \] To convert an angle i degrees to radian we must multiply the given angle by \[\huge \frac{\pi}{180^0}\]

OpenStudy (anonymous):

so what is theta in radians then?

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

@washcaps yes...

OpenStudy (anonymous):

I'm here

OpenStudy (anonymous):

I don't know what to put in the empty cell of "theta in radians"

OpenStudy (anonymous):

just post your question. Is this the above question?

OpenStudy (anonymous):

yes it is the same question. it all started with cos(theta) 1

OpenStudy (anonymous):

\[0^0 = 0^0 \times \frac {\pi}{180^0} = 0^c\]

OpenStudy (anonymous):

ok… so I put 0 degrees in the cell that says theta in degrees?

OpenStudy (anonymous):

for angle in0 radians there would be no change for the values of the table, means table would take the ame values.

OpenStudy (anonymous):

Medal please

OpenStudy (anonymous):

I will give you a medal. Just clarify my chart. It will be an attachment

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

just view my attachment and see if it is correct

OpenStudy (anonymous):

hey @surjithayer could you help?

OpenStudy (anonymous):

1. write 0,1,2,3,4 from left to right for sin and from right to left for cos 2.divide each by 4 3.take the square root of each. 4.for tan ,divide sin by cos